OpenStudy (bacon):

Anita and Winnie share some oranges and lemons in the ratio of 5:7. The oranges and lemons that Anita has are in the ratio of 3:5. If Anita has 30 fewer oranges than lemons, how many oranges and lemons does Winnie have? Is there an equation for this?

OpenStudy (anonymous):

Anita has a oranges and b lemons Winnie has c oranges and d lemons

OpenStudy (anonymous):

can u do it now?

OpenStudy (bacon):

sorry no, i don't really understand that

OpenStudy (anonymous):

a/c = 5/7 b/d = 5/7 a/b = 3/5 b - a = 30 now?